# The ambiguous AT2022rze: Changing-look AGN mimicking a supernova in a merging galaxy system

### v1

@@ Abstract
Core collapse supernovae (CCSNe) mark the explosive deaths of massive stars (> 8M⊙). CCSNe can be classified by spectroscopic and/or photometric features (e.g: Gal-Yam, 2017). In particular, events that display exceptionally high peak luminosities, typically brighter than 20 mag in optical bands, are classified as SLSNe (e.g: Gal-Yam, 2019). When hydrogen (H) lines are visible in the spectra of these events, they are classified as Type II SLSNe or Type IIn SLSNe if the hydrogen features are narrow, which indicates interaction between the SN ejecta and a dense circumstellar material (CSM, e.g: Smith, 2017). Tidal disruption events (TDEs) occur when a star intersects the tidal radius of a black hole (BH) and is torn apart by tidal forces, producing a luminous flare powered by the fallback and accretion of stellar debris (e.g., Gezari, 2021). Active galactic nuclei (AGNs) are sustained by accretion of gas onto supermassive BHs (SMBHs) at galactic centers, producing intense, persistent emission across the electromagnetic spectrum (e.g., Padovani et al., 2017). A subclass of AGNs have been discovered to change their aspect and/or ignite when their host galaxies transition from quiescence to active accretion into the central BH, displaying luminous, energetic light curves. These have been labeled changing-look AGN (CLAGN, e.g., Ricci & Trakhtenbrot, 2023).
@@
In most cases, distinguishing SLSNe, TDEs and AGNs is relatively straightforward by spectral matching or light curve analysis (a large variety of classification tools exits, here we only list a small fraction as example: Howell et al., 2005; Blondin & Tonry, 2007; Narayan et al., 2018; Muthukrishna et al., 2019; Förster et al., 2021; Fraga et al., 2024). For instance, AGN variability is generally stochastic and persistent, unlike the more defined temporal structure of SLSNe and TDEs. TDEs typically appear blue and exhibit little color evolution, whereas SLSNe begin with high temperatures and gradually redden as they cool. On the contrary, CLAGNs are very diverse, and their characteristics can resemble those of SLSNe IIn and TDEs (e.g., Komossa & Grupe, 2024).
@@
Recent studies have started to incorporate host galaxy properties alongside photometric behavior to aid in classification (e.g: Gomez et al., 2020; Gagliano et al., 2023). However, overlap in host galaxy location, particularly for events located in nuclear or circumnuclear regions, makes classification even more ambiguous (e.g., French et al., 2020; Ward et al., 2021; Grishin et al., 2021; Li et al., 2023). Additionally, dust and gas surrounding the transient may produce prolonged, bumpy, and luminous light curves, further obscuring classification. As a result SLSNe, TDEs, and (CL)AGNs can become contaminants in corresponding sample studies (see for example the contaminants in the SLSN II sample presented by Pessi et al., 2024).
@@
There is an increasing number of energetic transients with ambiguous characteristics being discovered (e.g. Kankare et al., 2017). A large fraction of these events is found at the nuclei of their host galaxies and have been dub these events Ambiguous Nuclear Transients (ANTs, Wiseman et al., 2025). ANTs exhibit heterogeneous spectral and photometric properties that do not really fit those of SLSNe, TDEs, or AGNs. Mid-infrared observations suggest that dust and molecular gas may shape the optical properties of ANTs, highlighting the need for multiwavelength observations to understand their nature. AT2022rze, exemplifies this ambiguity. It appears centrally located in a compact structure, possibly a small galaxy merging with a larger host, but exhibits properties that defy a straightforward classification.
@@ 2 Observations
The discovery of AT2022rze (also known as Gaia22dmp, ZTF22abnfjsm and ATLAS22bmgb) was reported on August 21st 2022 by Hodgkin et al. (2022). Its location at R.A.12h22m52.16s, Dec76∘02′47.26′′, (J2000.0) allows the identification of the galaxy WISEAJ122252.07760253.9 as the host. AT2022rze is located at the center of a small luminous “blob” (galaxy B, see Sect. 3.3) located South-East from the geometrical center of the host (see Fig. 1, see Sect. 3.3 for further discussion of the host). AT2022rze was classified as a SN IIn almost 4 months after its discovery, on December 1 2022, by Hinds et al. (2022). Two years later, on December 16 2024, it was re-classified as an AGN by (Gagliano, 2024). Both classifications are based on spectroscopic features, raising questions about the true nature of this event.
From the position of the hydrogen emission lines we derive a redshift of , consistent with the reported photometric redshift of the host (Bilicki et al., 2014). We employ the astropy.cosmology software adopting H km s-1 Mpc-1, , as cosmological parameters. With these, we obtain a distance modulus of mag. We do not consider uncertainties related to host galaxy peculiar velocities as it is in the Hubble flow () and corrections would be small. From the NASA/IPAC Extragalactic Database’s (NED111The NASA/IPAC Extragalactic Database (NED) is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.) we obtain a Milky Way (MW) extinction of A mag. We use NED’s Galactic Extinction Calculator222NED’s Extinction Calculator considers the recalibration presented by Schlafly & Finkbeiner (2011) to the extinction map presented by Schlegel et al. (1998), assuming a Fitzpatrick (1999) reddening law with R., accessed through the ned_extinction_calc script333https://github.com/mmechtley/ned_extinction_calc to obtain the MW extinction in different photometric bands. The effective wavelength corresponding to each filter was obtained from the Spanish Virtual Observatory Filter Information service (SVO, Rodrigo et al., 2012; Rodrigo & Solano, 2020). We do not consider host extinction, although see Sect. 3.3. The available photometric and spectroscopic data are presented below.
@@ 2.1 Photometry
AT2022rze was detected by the public surveys carried out with the Gaia satellite (Gaia Collaboration et al., 2016), the Asteroid Terrestrial-impact Last Alert System (ATLAS, Tonry et al., 2018; Smith et al., 2020) and the Zwicky Transient Facility (ZTF, Bellm et al., 2019a; Graham et al., 2019; Masci et al., 2019; Dekany et al., 2020). Thus, it has extensive optical photometric follow up. Additional photometric follow up was obtained as part of the ZTF collaboration with the Spectral Energy Distribution Machine (SEDM, Blagorodnova et al., 2018) on the 60-inch telescope (P60) at Palomar Observatory. Moreover, mid-Infrared (MIR) observations were obtained by the Wide-field Infrared Survey Explorer (WISE) instrument.
The observed photometry is listed in Tables 2, 3 and 4. The corresponding light curves are shown in Fig. 2. Absolute magnitudes are obtained as , where is apparent magnitude in the AB system, mag is the distance modulus (see Sect. 1), is the MW extinction in the considered band and is the cosmological term for the K-correction (Hogg et al., 2002) obtained as (as presented in Chen et al. 2023 and Pessi et al. 2024).
@@ 2.1.1 Gaia
The Gaia Science Alerts project (Hodgkin et al., 2021) is dedicated to search for transient events within Gaia data. Alerts for discovered transients are made public through Gaia alerts that provide information about the transient, including the average magnitude of the source across all CCD strips at each considered timestamp. We obtained the public light curve from the dedicated repository444https://gsaweb.ast.cam.ac.uk/alerts/. The estimated absolute magnitudes are presented as black rhomboids in the bottom panel of Fig. 2. Since the public light curves do not include associated error bars, we do not include error bars here.
@@ 2.1.2 ATLAS
The ATLAS survey scans the sky in two wide filters: (or “cyan” band, in the wavelength range 4200–6500 Å, roughly corresponding to the typical range), and (or “orange” band, in the wavelength range of 5600–8200 Å, roughly corresponding to the typical range); with a cadence of around two days (Smith et al., 2020; Shingles et al., 2021). We retrieved the ATLAS forced-photometry from the dedicated repository555https://fallingstar-data.com/forcedphot/, and processed it utilizing the suggested pipeline developed by Young (2020), with the intra-night stacking option. The and light curves are presented in flux and absolute magnitude in the top and bottom panels of Fig. 2, respectively.
@@ 2.1.3 ZTF
ZTF is a wide-field (47-square-degree field of view), high-cadence survey (minutes to days, with a three day average for the public survey, Bellm et al. 2019b), that covers the whole northern sky. ZTF forced point-spread function (PSF) photometry is processed and distributed by the Science Data System at IPAC666Formerly referred to as the Infrared Processing & Analysis Center (https://www.ipac.caltech.edu). (Masci et al., 2019). We requested the available forced photometry following the steps outlined in the ZTF forced photometry guideline777http://web.ipac.caltech.edu/staff/fmasci/ztf/forcedphot.pdf. We process the retrieved light curves in the same way as outlined by Pessi et al. (2024), removing points with bad quality flags and finding the zero flux baseline corresponding to each filter. The light curves are presented in flux and absolute magnitude in the top and bottom panels of Fig. 2, respectively.
We obtained further P60 photometry through the ZTF collaboration. Photometry is obtained through PSF fitting, calibrated with standard stars from the Sloan Digital Sky Survey (SDSS, Gunn, 1995). All the photometry is processed using the Fpipe software (Fremling et al., 2016) using SDSS reference images.
@@ 2.1.4 WISE
The field of AT2022rze was observed with WISE in the W1 and W2 bands (with central wavelengths of 3.4 and 4.6 m, respectively) during the NEOWISE-reactivation mission (NEOWISE-R; Mainzer et al., 2014). Single exposures in the W1 and W2 bands were collected from the NASA/IPAC Infrared Science Archive (IRSA) and coadded using the ICORE service (Masci, 2013). Reference images were created by coadding data from 2015, which were subtracted from the subsequent images. We perform aperture photometry on the subtracted images using a standard 8.25″radius aperture (and corresponding aperture corrections), and zero magnitude fluxes for the WISE W1 and W2 bands of Jy and Jy (Wright et al., 2010a).
@@ 2.1.5 Host
The host galaxy has been observed by the Wide-field Infrared Survey Explorer mission (WISE, Wright et al., 2010b), the Very Large Array Sky Survey (VLASS, Lacy et al., 2020; Gordon et al., 2021), and eROSITA (Predehl et al., 2021).
MIR observations of the host are available in the AllWISE Source Catalog888https://wise2.ipac.caltech.edu/docs/release/allwise/. AllWISE combines both the WISE cryogenic and NEOWISE (Mainzer et al., 2011) post-cryogenic survey phases, providing observations at 3.4, 4.6, 12, and 22 m, corresponding to the so called W1, W2, W3, and W4 filters, respectively. A coordinate search considering a 10 cone radius retrieves two sources, one consistent with the NED coordinates of the host, and the other offset 0.4W, 4.5S from that position (see Fig. 1, see Sect. 3.3 for further discussion). The magnitudes in each of the WISE filters are listed in Table 1.
@@
Radio observations of the host are available in the VLASS archive999http://cutouts.cirada.ca/. We retrieve quick look single epoch radio continuum images from the VLASS, centered on the position of AT2022rze. These images sample the radio continuum between 2-4 GHz, and were taken on September 15 2017, September 3 2020, and March 5 2023. We tentatively detect radio continuum emission (2.5) close to the center of the host galaxy in the epoch 1 and 3 images, but do not find emission at the position of AT2022rze. In order to increase the signal-to-noise ratio of the images, we stack the three epochs together, co-adding the images weighted by their rms (respectively 0.12 mJy.beam-1, 0.15 mJy.beam-1, and 0.11 mJy.beam-1). In the stacked image, we detect the 3 GHz radio continuum emission from the host galaxy with flux density mJy. We do not detect radio emission from AT2022rze, but derive a 3 upper limit mJy within the beam (3"), assuming radio continuum emission from the source is unresolved.
@@ 2.2 Spectroscopy
In addition to the photometric follow up, we obtained several epochs of spectroscopic observations using the SEDM spectrograph, the Spectrograph for the Rapid Acquisition of Transients (SPRAT, Piascik et al., 2014) at the 2-m Liverpool Telescope, the Low-Resolution Imaging Spectrograph (LRIS, Oke et al., 1995) at the 10-m Keck I telescope, the Alhambra Faint Object Spectrograph and Camera (ALFOSC) at the Nordic Optical Telescope, and the Double Spectrograph (DBSP, Oke & Gunn, 1982) at the Palomar 200-inch telescope. The SEDM spectra were reduced using the pysedm fully automated integral field spectrograph pipeline (Rigault et al., 2019; Kim et al., 2022). SPRAT spectra were reduced using an adaptation of the FRODOspec pipeline (Barnsley et al., 2012). LRIS spectra were reduced using the LPipe automated reduction pipeline (Perley, 2019). ALFOSC spectra were reduced using the PypeIt python spectroscopic data reduction pipeline (Prochaska et al., 2020a, b). DBSP spectra were reduced using the DBSP_DRP automated spectroscopic data reduction pipeline (Mandigo-Stoba et al., 2022). Follow-up observations were coordinated using the Fritz instance of SkyPortal (van der Walt et al., 2019; Coughlin et al., 2023). The log of spectroscopic observations can be found in Table 6. The obtained spectral evolution is shown in Fig. 3.
Most of the spectra were obtained with the slit aligned along the parallactic angle to minimize atmospheric dispersion effects. The ALFOSC spectra obtained at 422.2 and 465.5 days were observed with a slit rotation that allowed to include both the transient and the host and thus, obtain two spectral epochs of the latter. These are shown in Fig. 7.
The position of the transient was also observed by the Dark Energy Spectroscopic Instrument (DESI, DESI Collaboration et al., 2016a, b) on March 15th 2022, five months before the transient was discovered. The spectrum is included in their recent data release (DESI Collaboration et al., 2025) and we include it in Fig. 7.

### v2

@@ Abstract
Core collapse supernovae (CCSNe) mark the explosive deaths of massive stars (> 8M⊙). CCSNe can be classified by spectroscopic and/or photometric features (e.g: 2017hsn..book..195G). In particular, events that display exceptionally high peak luminosities, typically brighter than 20 mag in optical bands, are classified as SLSNe (e.g: 2019ARA&A..57..305G). When hydrogen (H) lines are visible in the spectra of these events, they are classified as Type II SLSNe or Type IIn SLSNe if the hydrogen features are narrow, which indicates interaction between the SN ejecta and a dense circumstellar material (CSM, e.g: 2017hsn..book..403S). Tidal disruption events (TDEs) occur when a star intersects the tidal radius of a black hole (BH) and is torn apart by tidal forces, producing a luminous flare powered by the fallback and accretion of stellar debris (e.g., 2021ARA&A..59...21G). Active galactic nuclei (AGNs) are sustained by accretion of gas onto supermassive BHs (SMBHs) at galactic centers, producing intense, persistent emission across the electromagnetic spectrum (e.g., 2017A&ARv..25....2P). A subclass of AGNs have been discovered to change their aspect and/or ignite when their host galaxies transition from quiescence to active accretion into the central BH, displaying luminous, energetic light curves. These have been labeled changing-look AGN (CLAGN, e.g., 2023NatAs...7.1282R).
@@
In most cases, distinguishing SLSNe, TDEs and AGNs is relatively straightforward by spectral matching or light curve analysis (a large variety of classification tools exits, here we only list a small fraction as example: 2005ApJ...634.1190H; 2007ApJ...666.1024B; 2018ApJS..236....9N; 2019ApJ...885...85M; 2021AJ....161..242F; 2024A&A...692A.208F). For instance, AGN variability is generally stochastic and persistent, unlike the more defined temporal structure of SLSNe and TDEs. TDEs typically appear blue and exhibit little color evolution, whereas SLSNe begin with high temperatures and gradually redden as they cool. On the contrary, CLAGNs are very diverse, and their characteristics can resemble those of SLSNe IIn and TDEs (e.g., 2024SerAJ.209....1K).
@@
Recent studies have started to incorporate host galaxy properties alongside photometric behavior to aid in classification (e.g: 2020ApJ...904...74G; 2023ApJ...954....6G). However, overlap in host galaxy location, particularly for events located in nuclear or circumnuclear regions, makes classification even more ambiguous (e.g., 2020SSRv..216...32F; 2021ApJ...913..102W; 2021MNRAS.507..156G; 2023ApJ...950..161L). Additionally, dust and gas surrounding the transient may produce prolonged, bumpy, and luminous light curves, further obscuring classification. As a result SLSNe, TDEs, and (CL)AGNs can become contaminants in corresponding sample studies (see for example the contaminants in the SLSN II sample presented by 2025A&A...695A.142P).
@@
There is an increasing number of energetic transients with ambiguous characteristics being discovered (e.g. 2017NatAs...1..865K). A large fraction of these events is found at the nuclei of their host galaxies and have been dub these events Ambiguous Nuclear Transients (ANTs, 2025MNRAS.537.2024W). ANTs exhibit heterogeneous spectral and photometric properties that do not really fit those of SLSNe, TDEs, or AGNs. Mid-infrared observations suggest that dust and molecular gas may shape the optical properties of ANTs, highlighting the need for multiwavelength observations to understand their nature. AT2022rze, exemplifies this ambiguity. It appears centrally located in a compact structure, possibly a small galaxy merging with a larger host, but exhibits properties that defy a straightforward classification.
@@ 2 Observations
The discovery of AT2022rze (also known as Gaia22dmp, ZTF22abnfjsm and ATLAS22bmgb) was reported on August 21st 2022 by 2022TNSTR2428....1H. Its location at R.A.12h22m52.16s, Dec76∘02′47.26′′, (J2000.0) allows the identification of the galaxy WISEAJ122252.07760253.9 as the host. AT2022rze is located at the center of a small luminous “blob” (galaxy B, see Sect. 3.3) located South-East from the geometrical center of the host (see Fig. 1, see Sect. 3.3 for further discussion of the host). AT2022rze was classified as a SN IIn almost 4 months after its discovery, on December 1 2022, by 2022TNSCR3514....1H. Two years later, on December 16 2024, it was re-classified as an AGN by (2024TNSCR4932....1G). Both classifications are based on spectroscopic features, raising questions about the true nature of this event.
From the position of the hydrogen emission lines we derive a redshift of , consistent with the reported photometric redshift of the host (2014ApJS..210....9B). We employ the astropy.cosmology software adopting H km s-1 Mpc-1, , as cosmological parameters. With these, we obtain a distance modulus of mag. We do not consider uncertainties related to host galaxy peculiar velocities as it is in the Hubble flow () and corrections would be small. From the NASA/IPAC Extragalactic Database’s (NED111The NASA/IPAC Extragalactic Database (NED) is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology.) we obtain a Milky Way (MW) extinction of A mag. We use NED’s Galactic Extinction Calculator222NED’s Extinction Calculator considers the recalibration presented by 2011ApJ...737..103S to the extinction map presented by 1998ApJ...500..525S, assuming a 1999PASP..111...63F reddening law with R., accessed through the ned_extinction_calc script333https://github.com/mmechtley/ned_extinction_calc to obtain the MW extinction in different photometric bands. The effective wavelength corresponding to each filter was obtained from the Spanish Virtual Observatory Filter Information service (SVO, 2012ivoa.rept.1015R; 2020sea..confE.182R). We do not consider host extinction, although see Sect. 3.3. The available photometric and spectroscopic data are presented below.
@@ 2.1 Photometry
AT2022rze was detected by the public surveys carried out with the Gaia satellite (2016A&A...595A...1G), the Asteroid Terrestrial-impact Last Alert System (ATLAS, 2018PASP..130f4505T; 2020PASP..132h5002S) and the Zwicky Transient Facility (ZTF, 2019PASP..131a8002B; 2019PASP..131g8001G; 2019PASP..131a8003M; 2020PASP..132c8001D). Thus, it has extensive optical photometric follow up. Additional photometric follow up was obtained as part of the ZTF collaboration with the Spectral Energy Distribution Machine (SEDM, 2018PASP..130c5003B) on the 60-inch telescope (P60) at Palomar Observatory. Moreover, mid-Infrared (MIR) observations were obtained by the Wide-field Infrared Survey Explorer (WISE) instrument.
The observed photometry is listed in Tables 2, 3 and 4. The corresponding light curves are shown in Fig. 2. Absolute magnitudes are obtained as , where is apparent magnitude in the AB system, mag is the distance modulus (see Sect. 1), is the MW extinction in the considered band and is the cosmological term for the K-correction (2002astro.ph.10394H) obtained as (as presented in 2023ApJ...943...41C and 2025A&A...695A.142P).
@@ 2.1.1 Gaia
The Gaia Science Alerts project (2021A&A...652A..76H) is dedicated to search for transient events within Gaia data. Alerts for discovered transients are made public through Gaia alerts that provide information about the transient, including the average magnitude of the source across all CCD strips at each considered timestamp. We obtained the public light curve from the dedicated repository444https://gsaweb.ast.cam.ac.uk/alerts/. The estimated absolute magnitudes are presented as black rhomboids in the bottom panel of Fig. 2. Since the public light curves do not include associated error bars, we do not include error bars here.
@@ 2.1.2 ATLAS
The ATLAS survey scans the sky in two wide filters: (or “cyan” band, in the wavelength range 4200–6500 Å, roughly corresponding to the typical range), and (or “orange” band, in the wavelength range of 5600–8200 Å, roughly corresponding to the typical range); with a cadence of around two days (2020PASP..132h5002S; 2021TNSAN...7....1S). We retrieved the ATLAS forced-photometry from the dedicated repository555https://fallingstar-data.com/forcedphot/, and processed it utilizing the suggested pipeline developed by Young_plot_atlas_fp, with the intra-night stacking option. The and light curves are presented in flux and absolute magnitude in the top and bottom panels of Fig. 2, respectively.
@@ 2.1.3 ZTF
ZTF is a wide-field (47-square-degree field of view), high-cadence survey (minutes to days, with a three day average for the public survey, 2019PASP..131f8003B), that covers the whole northern sky. ZTF forced point-spread function (PSF) photometry is processed and distributed by the Science Data System at IPAC666Formerly referred to as the Infrared Processing & Analysis Center (https://www.ipac.caltech.edu). (2019PASP..131a8003M). We requested the available forced photometry following the steps outlined in the ZTF forced photometry guideline777http://web.ipac.caltech.edu/staff/fmasci/ztf/forcedphot.pdf. We process the retrieved light curves in the same way as outlined by 2025A&A...695A.142P, removing points with bad quality flags and finding the zero flux baseline corresponding to each filter. The light curves are presented in flux and absolute magnitude in the top and bottom panels of Fig. 2, respectively.
We obtained further P60 photometry through the ZTF collaboration. Photometry is obtained through PSF fitting, calibrated with standard stars from the Sloan Digital Sky Survey (SDSS, 1995AAS...186.4405G). All the photometry is processed using the Fpipe software (2016A&A...593A..68F) using SDSS reference images. The P60 photometry is neither color-corrected nor adjusted for differences in filter response. Although we do not use these measurements in our analysis, we include them here for completeness.
@@ 2.1.4 WISE
The field of AT2022rze was observed with WISE in the W1 and W2 bands (with central wavelengths of 3.4 and 4.6 m, respectively) during the NEOWISE-reactivation mission (NEOWISE-R; mainzer2014). Single exposures in the W1 and W2 bands were collected from the NASA/IPAC Infrared Science Archive (IRSA) and coadded using the ICORE service (masci2013). Reference images were created by coadding data from 2015, which were subtracted from the subsequent images. We perform aperture photometry on the subtracted images using a standard 8.25″radius aperture (and corresponding aperture corrections), and zero magnitude fluxes for the WISE W1 and W2 bands of Jy and Jy (Wright2010).
@@ 2.1.5 Host
The host galaxy has been observed by the Wide-field Infrared Survey Explorer mission (WISE, 2010AJ....140.1868W), the Very Large Array Sky Survey (VLASS, 2020PASP..132c5001L; 2021ApJS..255...30G), and eROSITA (2021A&A...647A...1P). We note that the host is formed by two galaxies undergoing a merging event, namely galaxy A and galaxy B (see Fig. 1 and Sect. 3.3 for a detailed description). Throughout this work, we refer to the entire system as the host, specifying galaxy A or B when relevant.
MIR observations of the host are available in the AllWISE Source Catalog888https://wise2.ipac.caltech.edu/docs/release/allwise/. AllWISE combines both the WISE cryogenic and NEOWISE (2011ApJ...731...53M) post-cryogenic survey phases, providing observations at 3.4, 4.6, 12, and 22 m, corresponding to the so called W1, W2, W3, and W4 filters, respectively. A coordinate search considering a 10 cone radius retrieves two sources, one consistent with the NED coordinates of the host, and the other offset 0.4W, 4.5S from that position (see Fig. 1, see Sect. 3.3 for further discussion). The magnitudes in each of the WISE filters are listed in Table 1.
@@
Radio observations of the host are available in the VLASS archive999http://cutouts.cirada.ca/. We retrieve quick look single epoch radio continuum images from the VLASS, centered on the position of AT2022rze. These images sample the radio continuum between 2-4 GHz, and were taken on September 15 2017, September 3 2020, and March 5 2023. We tentatively detect radio continuum emission (2.5) close to the NED coordinates of the host in the epoch 1 and 3 images, but do not find emission at the position of AT2022rze. In order to increase the signal-to-noise ratio of the images, we stack the three epochs together, co-adding the images weighted by their rms (respectively 0.12 mJy.beam-1, 0.15 mJy.beam-1, and 0.11 mJy.beam-1). In the stacked image, we detect the 3 GHz radio continuum emission from the host with flux density mJy. We do not detect radio emission from AT2022rze, but derive a 3 upper limit mJy within the beam (3"), assuming radio continuum emission from the source is unresolved.
@@ 2.2 Spectroscopy
In addition to the photometric follow up, we obtained several epochs of spectroscopic observations using the SEDM spectrograph, the Spectrograph for the Rapid Acquisition of Transients (SPRAT, 2014SPIE.9147E..8HP) at the 2-m Liverpool Telescope, the Low-Resolution Imaging Spectrograph (LRIS, 1995PASP..107..375O) at the 10-m Keck I telescope, the Alhambra Faint Object Spectrograph and Camera (ALFOSC) at the Nordic Optical Telescope, and the Double Spectrograph (DBSP, 1982PASP...94..586O) at the Palomar 200-inch telescope. The SEDM spectra were reduced using the pysedm fully automated integral field spectrograph pipeline (2019A&A...627A.115R; 2022PASP..134b4505K). SPRAT spectra were reduced using an adaptation of the FRODOspec pipeline (2012AN....333..101B). LRIS spectra were reduced using the LPipe automated reduction pipeline (2019PASP..131h4503P). ALFOSC spectra were reduced using the PypeIt python spectroscopic data reduction pipeline (2020JOSS....5.2308P; pypeit:joss_pub). DBSP spectra were reduced using the DBSP_DRP automated spectroscopic data reduction pipeline (2022JOSS....7.3612M). Follow-up observations were coordinated using the Fritz instance of SkyPortal (2019JOSS....4.1247V; 2023ApJS..267...31C). The log of spectroscopic observations can be found in Table 6. The obtained spectral evolution is shown in Fig. 3.
Most of the spectra were obtained with the slit aligned along the parallactic angle to minimize atmospheric dispersion effects. The ALFOSC spectra obtained at 422.2 and 465.5 days were observed with a slit rotation that allowed to include both the transient (located at the center of galaxy B) and galaxy A (see Sect. 3.3) and thus, obtain two spectral epochs of the latter. These are shown in Fig. 7.
The position of the transient (i.e. galaxy B) was also observed by the Dark Energy Spectroscopic Instrument (DESI, 2016arXiv161100036D; 2016arXiv161100037D) on March 15th 2022, five months before the transient was discovered. The spectrum is included in their recent data release (2025arXiv250314745D) and we include it in Fig. 7.